50 ml of 1 M oxalic acid is shaked with 0.5 g wood charcoal. The final concentration of the solution after adsorption is 0.5 M. What is the amount of oxalic acid absorbed per gram of carbon?
    (A)    3.15 g
    (B)    3.45 g
    (C)    6.30 g
    (D)    None

Grade:upto college level

1 Answers

Radhika Batra
10 years ago

Amount of oxalic acid initially present

    =(1×50)/1000×63=3.15 g

    Amount of oxalic acid after adsorption

    =50/1000×63/2=3.15/2=1.575 g

    Amount absorbed = 3.15 – 1.575 = 1.575 g

    Amount absorbed per gram = 1.575 × 2

                        = 3.15 gm